# A quantum implementation of high-order power method for estimating geometric entanglement of pure states

### v1

@@ 1 Introduction
For a quantum algorithm to provide advantage over a classical alternative, entanglement [1] is a necessary ingredient. Not only the presence but the “degree” of entanglement in a quantum state is an important property for many applications [2], for example quantum information technologies [3], quantum teleportation and quantum communication [1] and quantum cryptography [4]. On the one hand, if there is not “enough” entanglement a quantum circuit can be efficiently simulated by classical devices [5]. On the other hand, some systems with “maximally” entangled states, such as stabilizer codes [6], admit efficient classical simulations [7, 8]. Also, highly entangled states are not useful as computational resources in the measurement based computing paradigm [9]. In the setting of parameterised quantum circuits, entanglement contributes to “barren plateaus” in the cost function landscape that make training a challenge [10, 11].
@@
Consequently, there has been an effort to define and quantify [2] entanglement with different methods [12] being favoured depending on the application. Some definitions, (for example, quantum mutual information or von Neumann entropy [1]) are measures of bipartite entanglement only and are computationally expensive for mixed states. Others, such as concurrence [13], have no unique definition for higher ($>2$) dimensional systems [14, 15]. In fact, the very concept of $n$-partite (multi-qubit) entanglement is still not clearly understood [16].
@@
The geometric measure of entanglement ($E_{G}$[17, 18, 19] is a popular multi-partite entanglement metric with a clear geometric interpretation, which naturally extends to the $n$-partite case and to mixed states [19]. Geometric entanglement is, among other applications, useful to define entanglement witnesses [19] which are used as a positive indicator of entanglement for a subset of quantum states.
@@
The first algorithm to compute $E_{G}$ used it (under the name of “the Groverian measure of entanglement”) to assess the probability of success for an initial state in Grover’s Algorithm [20, 21]. This method was rephrased in terms of eigenvalues and Singular Value Decomposition (SVD) and extended for mixed states [22]. A similar method was formulated in terms of Tucker decomposition [23] using the Higher Order Orthogonal Iteration (HOOI) algorithm [24]. A parallel strand of work [25, 26], based on the fundamental connection between geometric entanglement and tensor theory, introduced a series of algorithms to approximate $E_{G}$ [27, 28, 29]. These approaches can be considered as generalizations of the popular power methods (e.g. the Lanczos method and Arnoldi iterations) used for studying the eigenvalues of matrices (for example see [30]).
@@
With accessible quantum hardware, measuring entanglement on a physical device becomes a problem of interest. The algorithms mentioned above are all intended to be executed on classical machines. Naïvely, we would have to do full-state tomography to reconstruct the quantum state in a classical computer to calculate its entanglement. As the number of qubits increases the amount of measurements required for tomography rises exponentially. Alternatively, if we know the quantum state in advance we can prepare the appropriate entanglement witness [3]. However, the latter method does not allow one to get the measure of entanglement itself. Recently variational quantum circuits (VQC) have been suggested to compute geometric entanglement of pure states on a quantum computer [31, 32]. VQC algorithms, unfortunately, suffer from “barren plateaus” [10, 11] which hinders the scalability of the approach.
@@
In this paper, we present an iterative quantum algorithm for computing the geometric entanglement of pure quantum states. The algorithm is a quantum adaptation of the Higher Order Power Method (HOPM) [24] for estimating Rank-1 Tensor Approximation (RTA). We show how to implement crucial steps of HOPM in the quantum domain and analyse their robustness w.r.t. noise. Our aim is to measure entanglement on near-term quantum devices more time efficiently than full-state tomography and more space efficiently (qubits vs classical memory) than executing HOPM on a classical device.
@@ 2.1 Preliminaries: Geometric Entanglement in terms of rank-1 tensor approximation
$E_{G}(\ket{\psi})=1-\max_{\ket{\phi}\in S_{n}}|\langle\phi|\psi\rangle|^{2},$ (2)
where $\tilde{\lambda}=\max\limits_{\ket{\phi}\in S_{n}}|\langle\phi|\psi\rangle|$ | is called the entanglement eigenvalue.
One approach to compute the value of $\tilde{\lambda}$ is to solve the problem of finding the “closest” fully separable state (CSS) |$\ket{\phi}$ to the state |$\ket{\psi}$. CSS is formally stated as a minimization problem over |$\ket{\phi}\in S_{n}$ with the objective function being the distance between |$\ket{\psi}$ and |$\ket{\phi}$:
@@
where $d(a,b)=\|a-b\|_{F}$ is the reference distance with $a,b\in\mathbb{H}^{n}$ and $\|\cdot\|_{F}$ being the Frobenius norm (we will omit the subscript $F$ in what follows). We note that the minimizer of Eq. (3) exists (but may not be unique) since the set of fully separable states $S_{n}$ is the classical Segre variety [33, 34, 35], which is closed in Zariski and Euclidean topology when defined over the complex numbers. Minimizing the distance from a variety is a problem well studied from different perspectives, for example in the context of algebraic geometry the concept of Euclidean Distance Degree has been introduced [36]. In the present work we deal with the minimization of Eq. (3) in terms of a system of nonlinear equations with $\lambda$ being a Lagrange multiplier (see [26, 19, 29]):
$T_{\psi}\times_{\hat{i}}\left({\bf v}_{1}^{\ast},\ldots,{{\bf v}_{i-1}^{\ast}},{\bf v}_{i+1}^{\ast},\ldots,{\bf v}_{n}^{\ast}\right)=\lambda{\bf v}_{i},$ (4)
$T_{\psi}^{*}\times_{\hat{i}}\left({{\bf v}_{1}},\ldots,{{\bf v}_{i-1}},{{\bf v}_{i+1}},\ldots,{{\bf v}_{n}}\right)=\lambda{\bf v}_{i}^{\ast},$
$\|{\bf v}_{i}\|=\sum_{b_{i}}|v_{i,b_{i}}|^{2}=1,\quad i\in\{1,\ldots,n\}$
$\lambda\in\mathbb{R},$
@@
and ${\bf v}_{i}=(v_{i,0},v_{i,1})^{\rm T}\in\mathbb{C}_{2}$ is a vector of components of the one-qubit states |$\ket{{\bf v}_{i}}$, defined in Eq. (1), in the same basis; the asterisk means complex conjugate; the symbol ×$\times_{\hat{i}}$ is $n$-mode vector product over all modes except the $i$-th one, which is a contraction of a tensor on the left with the tuple of vectors on the right, skipping the $i$-th index (mode) of the tensor (when the subscript is not given none of the modes are skipped). In tensor analysis, the problem of finding a solution to a system of equations such as Eq. (4) is usually referred to as the $U$-eigenpair problem for the tensor $T_{\psi}$. One way to solve this problem is by using RTA [26, 29].
The problem of RTA for a particular tensor is $\mathsf{NP}$-hard [37] and so exact or global techniques such as homotopy continuation methods [38] may take exponential time or fail to converge. Alternatively, approximate techniques may run quickly but may return a local minima. Some of the well-known algorithms that can be used for approximating RTA are Higher-Order Singular Value Decomposition (HOSVD) [39], HOPM and HOOI [24]. In the particular case of tensors representing the pure quantum state of a finite system of qubits, both HOSVD and HOOI simplify to the HOPM algorithm, which is based upon the Alternating Least Squares (ALS) approach [24, 40] for solving non-linear systems of equations.
@@
1:Let $(\mathbf{v}_{1}^{(0)},\ldots,\mathbf{v}_{n}^{(0)})$ be a tuple of some $n$ one-qubit state vectors
@@
4:${\bf u}^{(k+1)}_{i}=T_{\psi}\times_{\hat{i}}\left({{\bf v}^{(k+1)^{\ast}}_{1}},\ldots,{{\bf v}^{(k+1)^{\ast}}_{i-1}},{\bf v}^{(k)^{\ast}}_{i+1},\ldots,{\bf v}^{(k)^{\ast}}_{n}\right)$
@@
8:$\lambda^{(k+1)}=\left|T_{\psi}\times\left({{\bf v}^{(k+1)^{\ast}}_{1}},\ldots,{\bf v}^{(k+1)^{\ast}}_{n}\right)\right|$
@@ 2.2 Quantum implementation of HOPM for estimating the entanglement eigenvalue
In this section we present an iterative quantum algorithm, which is a HOPM approach for approximating RTA with its crucial steps (lines 4-8 in Algorithm 1) performed on a quantum device. We will refer to this algorithm as QHOPM. The input for the algorithm is an $n$-qubit quantum circuit $U_{\psi}$ that prepares the target state |$\ket{\psi}=U_{\psi}\ket{\bf 0}\in\mathbb{H}^{n}$, where |$\ket{\bf 0}=\ket{0}^{\otimes n}$ is the initial state of the system. The separable state |$\ket{\phi}\in S_{n}$ is encoded as a tensor product of one-qubit $x$ and $z$ rotations acting on |$\ket{\bf 0}$:
@@
where |$\ket{\mathbf{v}_{i}}$ are one-qubit states and $\vartheta\in[0,\pi),\varphi\in[0,2\pi)$. This choice of encoding of |$\ket{\mathbf{v}_{i}}$ differs from any other encoding only by global phase and therefore is valid due to Remark 3. We choose an initial separable state |$\ket{\phi^{(0)}}$ by randomly choosing the angles $(\vartheta_{i}^{(0)},\varphi_{i}^{(0)})$ as a starting point for the approximation.
@@
Lines 4 and 6 of Algorithm 1 for the state |$\ket{\psi}=U_{\psi}\ket{\bf 0}$ at any iteration $k>0$ are equivalent, up to a global phase, to the following update of |$\ket{\mathbf{v}_{i}^{(k+1)}}$:
|$\ket{{\bf v}_{i}^{(k+1)}}=\sum_{b}v_{i,b}^{(k+1)}\ket{b}=\frac{\sum_{b}u_{i,b}^{(k+1)}\ket{b}}{\sqrt{\sum_{b}|u_{i,b}^{(k+1)}|^{2}}},$ (6)
@@
$u_{i,b}^{(k+1)}=\left(\bra{0}^{\otimes(n-i)}\otimes\bra{b}\otimes\bra{0}^{\otimes(i-1)}\right){V_{i}^{(k+1)}}^{\dagger}U_{\psi}\ket{\bf 0}$ (7)
@@
$\begin{split}V_{i}^{(k+1)}=&\left[\bigotimes_{l=i+1}^{n}R_{z}(\varphi_{l}^{(k+1)})R_{x}(\vartheta_{l}^{(k+1)})\right]\\ &\otimes I\otimes\\ &\left[\bigotimes_{l=1}^{i-1}R_{z}(\varphi_{l}^{(k)})R_{x}(\vartheta_{l}^{(k)})\right],\end{split}$ (8)
(see Fig. 1) and |$\ket{b}$ is a one-qubit computational basis.
@@
$u_{i,b}^{(k+1)}=T_{\psi}\times\left({\textbf{v}_{1}^{(k+1)}}^{\ast},\ldots,{\textbf{v}_{i-1}^{(k+1)}}^{\ast},{\bf e}^{\ast}_{b},{\textbf{v}_{i+1}^{(k)}}^{\ast},\ldots,{\textbf{v}_{n}^{(k)}}^{\ast}\right),$
where ${\bf e}_{b}$ is a one-qubit basis vector for the $b$-th component. Let us substitute to the right-hand side the following expressions for the components of $T_{\psi}$ and ${\bf v}_{m\neq i}^{(k)}$ in terms of $U_{\psi}$ and $R_{z}R_{x}$:
$\psi_{l_{1},...,l_{n}}=\bra{l_{1}...l_{n}}U_{\psi}\ket{\bf 0};\quad v_{m,l}^{(k)}=\bra{l}R_{z}(\varphi_{m}^{(k)})R_{x}(\vartheta_{m}^{(k)})\ket{0},$
and taking into account that the $l$-th component of ${\bf e}_{b}$ in bra-ket notation is $\bra{l}b\rangle$, we get for the $b$-th component of $\textbf{u}_{i}^{(k+1)}$
$\begin{split}u_{i,b}^{(k+1)}=&\sum_{l_{1},...,l_{n}}\bra{0}R_{x}^{\dagger}(\vartheta_{1}^{(k+1)})R_{z}^{\dagger}(\varphi_{1}^{(k+1)})\ket{l_{1}}...\langle b\ket{l_{i}}...\\ &\bra{0}R_{x}^{\dagger}(\vartheta_{n}^{(k)})R_{z}^{\dagger}(\varphi_{n}^{(k)})\ket{l_{n}}\bra{l_{1}...l_{n}}U_{\psi}\ket{0}\\ =&\left(\bra{0}^{\otimes(n-i)}\otimes\bra{b}\otimes\bra{0}^{\otimes(i-1)}\right){V_{i}^{(k+1)}}^{\dagger}U_{\psi}\ket{\bf 0},\end{split}$
where $V_{i}^{(k+1)}$ is defined by Eq. (8). This expression corresponds to Eq. (7).
@@
Proposition 4 allows us to understand which type of operations and measurements we should perform to implement lines  4 and 6 of the algorithm, in particular, we need to be able to measure $u_{i,b}^{(k+1)}$ and obtain the updated pair of angles $(\vartheta_{i}^{(k+1)},\varphi_{i}^{(k+1)})$ for further encoding the updated |$\ket{{\bf v}_{i}^{(k+1)}}$.

### v2

@@ 1 Introduction
For a quantum algorithm to have an advantage over a classical alternative, entanglement [1] is a necessary ingredient. Not only the presence, but the “degree” of entanglement in a quantum state is an important property for many applications [2], for example quantum information technologies [3], quantum teleportation and quantum communication [1] and quantum cryptography [4]. On the one hand, if there is not “enough” entanglement, a quantum circuit can be efficiently simulated by classical devices [5]. On the other hand, some systems with “maximally” entangled states, such as stabiliser codes [6], admit efficient classical simulations [7, 8]. In addition, highly entangled states are not useful as computational resources in the measurement-based computing paradigm [9]. In the setting of parameterised quantum circuits, entanglement contributes to “barren plateaus” in the cost function landscape that make training a challenge [10, 11].
@@
Consequently, there has been an effort to define and quantify [2] entanglement, with different methods [12] being preferred depending on the application. Some definitions (for example, quantum mutual information or von Neumann entropy [1]) are measures of bipartite entanglement only and are computationally expensive for mixed states. Others, such as concurrence [13], have no unique definition for higher ($>2$) dimensional systems [14, 15]. In fact, the very concept of $n$-partite (multi-qubit) entanglement is still not clearly understood [16].
@@
The geometric measure of entanglement ($E_{G}$[17, 18, 19] is a popular multipartite entanglement metric with a clear geometric interpretation, which naturally extends to the $n$-partite case and to mixed states [19]. Geometric entanglement is, among other applications, useful when defining entanglement witnesses [19] which are used as a positive indicator of entanglement for a subset of quantum states.
@@
The first algorithm to compute $E_{G}$ used it (under the name of “the Groverian measure of entanglement”) to assess the probability of success for an initial state in Grover’s algorithm [20, 21]. This method was rephrased in terms of eigenvalues and Singular Value Decomposition (SVD) and extended to mixed states [22]. A similar method was formulated in terms of Tucker decomposition [23] using the Higher Order Orthogonal Iteration (HOOI) algorithm [24]. A parallel strand of work [25, 26], based on the fundamental connection between geometric entanglement and tensor theory, introduced a series of algorithms to approximate $E_{G}$ [27, 28, 29]. These approaches can be considered as generalisations of the popular power methods (e.g. the Lanczos method and Arnoldi iterations) used to study the eigenvalues of matrices (for example, see [30]).
@@
With accessible quantum hardware, measuring entanglement on a physical device becomes a problem of interest. The algorithms mentioned above are all intended to be executed on classical machines. Naïvely, we would have to do full-state tomography to reconstruct the quantum state in a classical computer to calculate its entanglement. As the number of qubits increases, the number of measurements required for tomography increases exponentially. Alternatively, if we know the quantum state in advance, we can prepare the appropriate entanglement witness [3]. However, the latter method does not allow one to obtain the measure of entanglement itself. Recently, variational quantum circuits (VQC) have been suggested to compute geometric entanglement of pure states on a quantum computer [31, 32]. VQC algorithms, unfortunately, suffer from “barren plateaus” [10, 11] which hinders the scalability of the approach.
@@
In this paper, we present an iterative quantum algorithm for computing the geometric entanglement of pure quantum states. The algorithm is a quantum adaptation of the Higher Order Power Method (HOPM) [24] to find solutions for Rank-1 Tensor Approximation (RTA). We show how to implement crucial steps of HOPM in the quantum domain and analyse their robustness w.r.t. noise. Our aim is to measure entanglement on near-term quantum devices more time efficiently than full-state tomography and more space efficiently (qubits versus classical memory) than executing HOPM on a classical device.
@@ 2.1 Preliminaries: Geometric Entanglement in terms of rank-1 tensor approximation
$\hat{E}_{G}(\ket{\psi})=1-\max_{\ket{\phi}\in S_{n}}|\langle\phi|\psi\rangle|^{2},$ (2)
where $\hat{\lambda}=\max\limits_{\ket{\phi}\in S_{n}}|\langle\phi|\psi\rangle|$ | is called the entanglement eigenvalue.
One approach to compute the value of $\hat{\lambda}$ is to solve the problem of finding the “closest” fully separable state |$\ket{\phi}$ to the state |$\ket{\psi}$. The closest separable state is formally stated as a minimization problem over |$\ket{\phi}\in S_{n}$ with the objective function being the distance between |$\ket{\psi}$ and |$\ket{\phi}$:
@@
where $d(a,b)=\|a-b\|_{F}$ is the reference distance with $a,b\in\mathbb{H}^{n}$ and $\|\cdot\|_{F}$ being the Frobenius norm (we will omit the subscript $F$ in what follows). We note that the minimiser of Eq.  (3) exists (but may not be unique) since the set of fully separable states $S_{n}$ is the classical Segre variety [33, 34, 35], which is closed in Zariski and Euclidean topology when defined over the complex numbers. Minimizing the distance from a variety is a problem well studied from different perspectives, for example in the context of algebraic geometry the concept of Euclidean Distance Degree has been introduced [36]. In the present work we deal with the minimization of Eq. (3) in terms of a system of nonlinear equations with $\lambda$ being a Lagrange multiplier (see [26, 19, 29]):
$T_{\psi}\times_{{i}}\left({\bf v}_{1}^{\ast},\ldots,{{\bf v}_{i-1}^{\ast}},{\bf v}_{i+1}^{\ast},\ldots,{\bf v}_{n}^{\ast}\right)=\lambda{\bf v}_{i},$ (4)
$T_{\psi}^{*}\times_{{i}}\left({{\bf v}_{1}},\ldots,{{\bf v}_{i-1}},{{\bf v}_{i+1}},\ldots,{{\bf v}_{n}}\right)=\lambda{\bf v}_{i}^{\ast},$
$\|{\bf v}_{i}\|^{2}=\sum_{b_{i}}|v_{i,b_{i}}|^{2}=1,\quad i\in\{1,\ldots,n\},\quad\lambda\in\mathbb{R},$
@@
and ${\bf v}_{i}=(v_{i,0},v_{i,1})^{\rm T}\in\mathbb{C}_{2}$ is a vector of components of the one-qubit states |$\ket{{\bf v}_{i}}$, defined in Eq.  (1), in the same basis; the asterisk means complex conjugate; the symbol ×$\times_{i}$ is $n$-mode vector product over all modes except the $i$-th one, which is a contraction of a tensor on the left with the tuple of vectors on the right, skipping the $i$-th index (mode) of the tensor (when the subscript is not given none of the modes are skipped). In tensor analysis, the problem of finding a solution to a system of equations such as Eq.  (4) is usually referred to as the $U$-eigenpair problem for the tensor $T_{\psi}$. One way to solve this problem is by reducing it to RTA [26, 29].
The problem of RTA is $\mathsf{NP}$-hard [37] and so exact or global techniques such as homotopy continuation methods [38] may take exponential time. Alternatively, approximate techniques may run quickly, but may return a local minima. Some of the well-known algorithms that can be used for approximating RTA are Higher-Order Singular Value Decomposition (HOSVD) [39], HOPM and HOOI [24]. In the particular case of tensors representing the pure quantum state of a finite system of qubits, both HOSVD and HOOI simplify to the HOPM algorithm, which is based upon the Alternating Least Squares (ALS) approach [24, 40] for solving non-linear systems of equations.
@@
1:Let $(\mathbf{v}_{1}^{(0)},\ldots,\mathbf{v}_{n}^{(0)})$, $\lambda^{(0)}$ be a tuple of some $n$ one-qubit state vectors and the corresponding entanglement eigenvalue.
@@
4:${\bf u}^{(k)}_{i}=T_{\psi}\times_{{i}}\left({{\bf v}^{(k)^{\ast}}_{1}},\ldots,{{\bf v}^{(k)^{\ast}}_{i-1}},{\bf v}^{(k-1)^{\ast}}_{i+1},\ldots,{\bf v}^{(k-1)^{\ast}}_{n}\right)$
@@
8:$\lambda^{(k)}=\left|T_{\psi}\times\left({{\bf v}^{(k)^{\ast}}_{1}},\ldots,{\bf v}^{(k)^{\ast}}_{n}\right)\right|$
@@ 2.2 Quantum implementation of HOPM for estimating the entanglement eigenvalue
In this section, we present an iterative quantum algorithm (Algorithm 2), which is a HOPM approach for approximating RTA with its crucial steps (lines 4-8 in Algorithm 1) performed on a quantum device. We will refer to this algorithm as QHOPM. The input for the algorithm is an $n$-qubit unitary operator $U_{\psi}\in\mathrm{U}(2^{n})$ that prepares the target state |$\ket{\psi}=U_{\psi}\ket{\bf 0}\in\mathbb{H}^{n}$, where |$\ket{\bf 0}=\ket{0}^{\otimes n}$ is the initial state of the system. The separable state |$\ket{\phi}\in S_{n}$ is encoded as a tensor product of one-qubit $x$ and $z$ rotations acting on |$\ket{\bf 0}$:
@@
where |$\ket{\mathbf{v}_{i}}$ are one-qubit states, and $\vartheta_{i}\in[0,\pi),\varphi_{i}\in[0,2\pi)$ are the angles used to encode |$\ket{\phi}$. This choice of encoding of |$\ket{\mathbf{v}_{i}}$ differs from any other encoding only by global phase and therefore is valid due to Remark 3. We choose an initial separable state |$\ket{\phi^{(0)}}$ by randomly choosing the angles $(\vartheta_{i}^{(0)},\varphi_{i}^{(0)})$ as a starting point for the approximation.
@@
Lines 4 and 6 of Algorithm 1 for the state |$\ket{\psi}=U_{\psi}\ket{\bf 0}$ at any iteration $k>0$ are equivalent, up to a global phase, to the following update of |$\ket{\mathbf{v}_{i}^{(k)}}$:
|$\ket{{\bf v}_{i}^{(k)}}=\sum_{b}v_{i,b}^{(k)}\ket{b}=\frac{\sum_{b}u_{i,b}^{(k)}\ket{b}}{\sqrt{\sum_{b}|u_{i,b}^{(k)}|^{2}}},$ (6)
@@
$u_{i,b}^{(k)}=\bra{b_{[i]}}{V_{i}^{(k)}}^{\dagger}U_{\psi}\ket{\bf 0},\quad\ket{b_{[i]}}=\left(\ket{0}^{\otimes(n-i)}\otimes\ket{b}\otimes\ket{0}^{\otimes(i-1)}\right)$ (7)
@@
$\begin{split}V_{i}^{(k)}=&\left[\bigotimes_{l=i+1}^{n}R_{z}(\varphi_{l}^{(k)})R_{x}(\vartheta_{l}^{(k)})\right]\\ &\otimes I\otimes\\ &\left[\bigotimes_{l=1}^{i-1}R_{z}(\varphi_{l}^{(k-1)})R_{x}(\vartheta_{l}^{(k-1)})\right],\end{split}$ (8)
(see Fig. 1), and |$\ket{b}$ is a one-qubit computational basis.
@@
$u_{i,b}^{(k)}=T_{\psi}\times\left({\textbf{v}_{1}^{(k)}}^{\ast},\ldots,{\textbf{v}_{i-1}^{(k)}}^{\ast},{\bf e}^{\ast}_{b},{\textbf{v}_{i+1}^{(k-1)}}^{\ast},\ldots,{\textbf{v}_{n}^{(k-1)}}^{\ast}\right),$
where ${\bf e}_{b}$ is a one-qubit basis vector for the $b$-th component. Let us substitute to the right-hand side the following expressions for the components of $T_{\psi}$ and ${\bf v}_{m\neq i}^{(t)}$ in terms of $U_{\psi}$ and $R_{z}R_{x}$:
$\psi_{l_{1},...,l_{n}}=\bra{l_{1}...l_{n}}U_{\psi}\ket{\bf 0};\quad v_{m,l}^{(t)}=\bra{l}R_{z}(\varphi_{m}^{(t)})R_{x}(\vartheta_{m}^{(t)})\ket{0},$
and taking into account that the $l$-th component of ${\bf e}_{b}$ in bra-ket notation is $\bra{l}b\rangle$, we get for the $b$-th component of $\textbf{u}_{i}^{(k)}$
$\begin{split}u_{i,b}^{(k)}=&\sum_{l_{1},...,l_{n}}\bra{0}R_{x}^{\dagger}(\vartheta_{1}^{(k)})R_{z}^{\dagger}(\varphi_{1}^{(k)})\ket{l_{1}}...\langle b\ket{l_{i}}...\\ &\bra{0}R_{x}^{\dagger}(\vartheta_{n}^{(k-1)})R_{z}^{\dagger}(\varphi_{n}^{(k-1)})\ket{l_{n}}\bra{l_{1}...l_{n}}U_{\psi}\ket{0}\\ =&\bra{b_{[i]}}{V_{i}^{(k)}}^{\dagger}U_{\psi}\ket{\bf 0},\end{split}$
where $V_{i}^{(k)}$ is defined by Eq.  (8). This expression corresponds to Eq. (7).
@@
Proposition 4 allows us to understand which type of operations and measurements we should perform to implement lines  4 and 6 of the algorithm, in particular, we need to be able to obtain the updated pair of angles $(\vartheta_{i}^{(k)},\varphi_{i}^{(k)})$ for further encoding the updated |$\ket{{\bf v}_{i}^{(k)}}=R_{z}(\varphi_{i}^{(k)})R_{x}(\vartheta_{i}^{(k)})\ket{0}$.